# LLM-Select: Feature Selection with Large Language Models

### v1

@@ 4.1 Evaluation on Small-Scale Datasets
In Figure 3(a), we show the feature selection paths for LLM-Score, LLM-Rank, and LLM-Seq based on GPT-4, which overlap significantly on all datasets except Diabetes Progression and NYC Rideshare*. Figures 2 and 3(a) together illustrate that GPT-4 achieves consistently strong feature selection performance regardless of the selection mechanism; even the simple strategy of querying for an importance score one-feature-at-a-time can be just as effective as those that account for the other features. In Figure 3(b), we overall observe a similar trend for GPT-3.5, albeit less pronounced. Meanwhile, Figures A4–A6 in Appendix A.3.2 show that for the smaller Llama-2 models, the performance of LLM-Score, LLM-Rank, and LLM-Seq is less consistent across datasets, suggesting that the effectiveness of each method is increasingly variable across datasets as model scale decreases.
Result 3: Zero-shot prompting with no dataset-specific context and greedy decoding leads to strong performance (Figure 4).
To assess the sensitivity of LLM-Score to prompt design and decoding strategy variations, we evaluate the downstream test performance across six prompt designs (described earlier in Section 4) and two decoding strategies (greedy/self-consistency). We use the change in the area under the feature selection paths to measure the impact of each variation. For classification tasks, an increase in the area indicates improved performance, as it suggests that fewer features are needed to achieve high AUROC. Conversely, for regression tasks, a decrease in the area indicates improvement. Thus, we quantify improvement in feature selection performance by computing the % increase for the classification tasks and the % decrease for the regression tasks. Figure 4 shows the average improvement for each (prompt design, decoding strategy) pair—with respect to that achieved with the default prompt and greedy decoding (). For the settings where we add examples, we only consider a one-shot setting given the relatively small number of features in the datasets. We find that no prompting strategy consistently improves feature selection performance across all LLMs, sometimes even degrading it (e.g., GPT-3.5, Llama-2 (13B)). Meanwhile, we find that the largest GPT-4 is substantially less sensitive to the prompting strategy and generally benefits from additional context, albeit to a limited extent. These observations suggest that zero-shot prompting without dataset-specific context, combined with greedy decoding, can be a sufficiently strong baseline prompting strategy for LLM-based feature selection.
Result 4: LLM-Score exhibits higher correlation with standard feature importance metrics as model scale increases (Figure 5).
@@
1.
SHAP (Lundberg & Lee, 2017),
2.
Fisher score (Duda et al., 2001),
3.
mutual information,
4.
Pearson correlation,
5.
Spearman correlation,
6.
permutation importance (Breiman, 2001).
We compute the Kendall’s coefficient (Kendall, 1938) to quantify the agreement in pairwise orderings, where +1/-1 indicates perfect agreement/disagreement. For SHAP, we compute the mean absolute Shapley value for each feature using all test samples, after training an XGBoost model (Chen & Guestrin, 2016). For permutation importance, we measure the average drop in test performance of an -penalized logistic/linear regression model when randomly shuffling the values of each feature 30 times. In Figure 5, we observe that there is no specific notion of importance that LLM-Score consistently aligns to. Meanwhile, as model scale increases, the LLM-generated scores generally exhibit higher correlation with the importance metrics considered (e.g., SHAP, Spearman). An interesting future research direction may be to investigate prompting strategies that increase alignment of LLM-Score to desired notions of importance. We include additional results on investigating the semantics of LLM-Score in Appendix A.3.6.
@@ 4.2 Evaluation on Large-Scale Datasets
We show that LLM-based feature selection also achieves strong performance on the more complex large-scale, higher dimensional datasets, each with 3000 features after preprocessing. Here, we focus on GPT-4-based LLM-Score with the default prompting setup, given its consistently strong performance on small-scale datasets and implementation efficiency. We construct supersets of the Income, Employment, Public Coverage, and Mobility datasets from folktables (Ding et al., 2021) by extracting all features available from the 2018 American Community Survey data for California while removing features that lead to label leakage. We also manually extract 3 datasets from the MIMIC-IV database (Johnson et al., 2023) for classifying whether an ICU patient was diagnosed with chronic kidney disease (CKD), chronic obstructive pulmonary disease (COPD), and heart failure (HF). We clarify that our MIMIC-IV datasets are not based on any existing data processing pipelines or splits previously used for MIMIC-IV. We provide full details in Appendix B.1.
We evaluate each feature selection method by measuring how the test performance of a downstream prediction model changes when we select the top 10%, 30%, 50%, 70% or 100% of all input concepts. We train the following models for downstream prediction: LightGBM (Ke et al., 2017), multi-layer perceptron (MLP), and -penalized logistic regression. For each method and dataset, we perform model selection via a random search with 40 hyperparameter samples. We average the test AUROC over 5 random seeds, which control the train-validation splits, the hyperparameter samples, and the initialization of model parameters.
@@
group LASSO (gLASSO; Yuan & Lin, 2006),
@@
MRMR (Ding & Peng, 2005),
@@
random feature selection,
@@
an all-features baseline.
Except for MRMR, we set up each method to select features at the concept level (e.g., via group-wise regularization), to match the behavior of LLM-Score. For LassoNet and gLASSO, we compute the regularization paths using warm starts (Friedman et al., 2010) (see Figure B1), and use the setting that selects about the same number of features as LLM-Score. For MRMR, we select exactly the same number of features as LLM-Score. We provide the remaining details in Appendix B.2.
@@
Figure 6(a) shows the downstream test AUROC, averaged over all of the folktables (left) and MIMIC-IV (right) datasets and all downstream prediction models, when selecting the top 30% of features according to each baseline and LLM-Score based on GPT-4. Here, we show the performance when selecting the top 30% of features, as the test performance roughly saturates to that of using all features when selecting more than 30% of features for all feature selection methods. We observe that GPT-4-based LLM-Score overall performs the best on the folktables datasets and significantly better than LassoNet and the random feature selection baseline on the MIMIC-IV datasets. In Figures 6(b–c), we show the full feature selection paths for the best-performing baseline (MRMR), LLM-Score, and random feature selection on the Income and Employment datasets from folktables and the CKD and COPD datasets from MIMIC-IV, with LightGBM as the downstream prediction model. We observe that for most of these datasets, only a small subset of features are highly predictive of the target outcome, as indicated by the high test AUROC of the best-performing baseline and the low test AUROC of the random feature selection baseline at the 10% and 30% marks. We find that LLM-Score is effective at selecting these highly predictive features, with performance rivaling that of the best-performing data-driven baseline and substantially better than random when selecting the top 10% and 30% of features. Notably, the strong performance on the manually extracted MIMIC-IV datasets suggests that LLM-Score can be effective even in complex domains like healthcare where, without access to the training data, substantial domain knowledge is required for feature selection (see Section B.1.2). We include the full results for all downstream prediction models on all datasets in Figure B2 in Appendix B.3.
@@ 5 Discussion and Conclusion
In this work, we demonstrated that LLMs are capable of performing feature selection for supervised learning tasks, even without access to the downstream training data. We showed that with sufficient LLM scale, selecting features by zero-shot prompting an LLM leads to strong downstream predictive performance, competitive with standard data-driven methods such as the LASSO. For the latest models such as GPT-4, we found that LLM-based feature selection consistently achieves strong performance across all selection strategies considered. Our findings suggest that LLMs may be useful not only for selecting the best features post data collection, but also for deciding what features to collect in the first place. This could potentially benefit practitioners in domains like healthcare where collecting high-quality data is costly.
Meanwhile, we note that LLMs may exhibit undesirable biases inherited from their pretraining data (Gallegos et al., 2024), potentially leading to biased feature selection and downstream performance disparities across data subpopulations. In the data-driven setting, it is possible to mitigate such issues by selecting features for each subpopulation independently or modifying the training objective to account for group fairness (Zemel et al., 2013; Sagawa et al., 2020; Izmailov et al., 2022). However, it is not immediately obvious how to incorporate similar notions of group fairness into LLM-based feature selection methods, especially for closed-source LLMs which can only be accessed via prompting. As such, combining LLM-driven feature selection with data-driven methods or using it in a human-in-the-loop setup may be a more reliable approach for mitigating bias concerns, especially in safety-critical domains. We leave an in-depth investigation of methods for incorporating group fairness to LLM-based feature selection for future work.
@@ Acknowledgements
We gratefully acknowledge DARPA (FA8750-23-2-1015), ONR (N00014-23-1-2368), NSF (IIS2211955), UPMC, Highmark Health, Abridge, Ford Research, Mozilla, the PwC Center, Amazon AI, JP Morgan Chase, the Block Center, the Center for Machine Learning and Health, and the CMU Software Engineering Institute (SEI) via Department of Defense contract FA8702-15-D-0002, for their generous support of our research. We also thank Michael Oberst, Goutham Rajendran, Bingbin Liu, and Che-Ping Tsai for helpful feedback and comments on this work.
@@ A.1 Datasets
For each dataset, we randomly shuffle and take a 80-20 train-test split. We then take a 5-fold split of the training set for cross-validation, where the 5-fold splits vary across the random seeds (=[1,2,3,4,5]) used throughout the experiments. The test set remains fixed and does not vary with the random seed used. For classification datasets, we always take a stratified split to preserve the label proportions across the train, validation, and test sets. We standardize all of the numerical features to have zero mean and unit variance, and one-hot encode all categorical features. Below, we provide the remaining details for each dataset.
@@
Credit-G (Hofmann, 1994) is a UCI dataset444https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data, where the goal is to predict whether a client at a bank carries high credit risk, given a set of attributes about the client (e.g., credit history, savings account status). The dataset contains 1000 samples and 61 features after preprocessing (20 concepts), with 700 positive samples and 300 negative samples.
@@
Bank (Moro et al., 2012) is a UCI dataset555https://archive.ics.uci.edu/dataset/222/bank+marketing, where the goal is to predict whether a client at a bank will subscribe to a term deposit, given data collected from a telemarketing campaign at a Portuguese banking institution from 2008 to 2013. The dataset contains 45211 samples and 51 features after preprocessing (16 concepts), with 5289 positive samples and 39922 negative samples.
@@
Give Me Some Credit is a Kaggle dataset666https://www.kaggle.com/c/GiveMeSomeCredit, where the goal is to predict whether an individual is likely to experience significant financial distress/delinquency within the next two years, given information about the individual’s financial status (e.g., debt ratio, monthly income). The dataset contains 120269 samples and 10 features after preprocessing (10 concepts), with 8357 positive samples and 111912 negative samples.
@@
COMPAS Recidivism is a dataset collected from a 2016 study777https://www.propublica.org/article/how-we-analyzed-the-compas-recidivism-algorithm on the racial biases present in the Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) algorithm (Larson et al., 2016). The goal is to predict whether a criminal defendant carries high risk of recidivism, given their criminal history and demographic attributes. The dataset contains 6172 samples and 23 features after preprocessing (14 concepts), with 2751 positive samples and 3421 negative samples.
@@
Pima Indians Diabetes (Smith et al., 1988) is a Kaggle dataset888https://www.kaggle.com/datasets/uciml/pima-indians-diabetes-database, where the goal is to predict whether a female adult patient of Pima Indian heritage has diabetes, given a set of clinical measurements and demographics. The dataset contains 768 samples and 8 features after preprocessing (8 concepts), with 268 positive samples and 500 negative samples.
@@
AUS Cars* is a Kaggle dataset999https://www.kaggle.com/datasets/nelgiriyewithana/australian-vehicle-prices published in 2023, where the goal is to predict whether the price of a car in Australia is above $30000, given a set of attributes about the car (e.g., year of manufacture, number of doors in the car, fuel consumption rate). The dataset contains 14283 samples and 33 features after preprocessing (11 concepts), with 6396 positive samples and 7887 negative samples.
@@
YouTube* is a Kaggle dataset101010https://www.kaggle.com/datasets/nelgiriyewithana/global-youtube-statistics-2023 published in 2023, where the goal is to predict whether YouTube channel has more than 20 million subscribers, given a set of attributes about the channel (e.g., total number of videos uploaded on the channel, date when the channel was created). The dataset contains 588 samples and 72 features after preprocessing (22 concepts), with 268 positive samples and 320 negative samples.
@@
CA Housing (Pace & Barry, 1997) is a StatLib dataset111111https://www.dcc.fc.up.pt/~ltorgo/Regression/cal_housing.html, where the goal is to predict the median housing price of a US Census block group in California, given data collected from the 1990 US Census. A block group, typically a population of 600 to 3000 people, is the smallest geographical unit for which the US Census Bureau publishes sample data. The dataset contains 20640 samples and 8 features after preprocessing (8 concepts).
@@
Diabetes Progression is a dataset121212https://www4.stat.ncsu.edu/~boos/var.select/diabetes.html used in a study by Efron et al. (2004), where the goal is to predict the disease progression level in diabetic patients, given their baseline blood serum measurements from the previous year and demographic information. The dataset contains 442 samples and 10 features after preprocessing (10 concepts).
@@
Wine Quality (Cortez et al., 2009) is a UCI dataset131313https://archive.ics.uci.edu/dataset/186/wine+quality collected from red and white vinho verde wine samples from northern Portugal, where the goal is to predict whether a wine is high or low quality, given its various physicochemical measurements (e.g., acidity, density). The dataset contains 6497 samples and 11 features after preprocessing (11 concepts).
@@
Miami Housing (Grinsztajn et al., 2022) is an OpenML dataset141414https://www.openml.org/search?type=data&sort=runs&id=43093&status=active, where the goal is to predict the selling price of a house in Miami in 2016, given structural (e.g., area, structure quality) and geographic (e.g., longitude, latitude) information about each house. The dataset contains 13932 samples and 15 features after preprocessing (15 concepts).
@@
Used Cars is a Kaggle dataset151515https://www.kaggle.com/datasets/vijayaadithyanvg/car-price-predictionused-cars, where the goal is to predict the selling price of a used car, given a set of attributes about each car (e.g., age, fuel type, number of previous owners). The dataset contains 301 samples and 11 features after preprocessing (7 concepts).
@@
NBA* is a Kaggle dataset161616https://www.kaggle.com/datasets/bryanchungweather/nba-player-stats-dataset-for-the-2023-2024 published in 2023, where the goal is to predict the number of points per game for an NBA basketball player, given a set of attributes and statistics (from the 2023–2024 NBA season) about the player (e.g., blocks per game, position, team). The dataset contains 388 samples and 61 features after preprocessing (27 concepts).
@@
NYC Rideshare* is a Kaggle dataset171717https://www.kaggle.com/datasets/aaronweymouth/nyc-rideshare-raw-data published in 2023, where the goal is to predict the total pay given to a rideshare-app driver in NYC after a trip, given information about the trip (e.g., time elapsed from ride request to dropoff, total distance of the trip). The original dataset on Kaggle contains some data from 2022, so we only extract the data rideshare data from 2023 for preprocessing and evaluation. The dataset contains 5000 samples and 19 features after preprocessing (15 concepts).

### v2

@@ 4.1 Evaluation on Small-Scale Datasets
The feature selection paths for LLM-Score, LLM-Rank, and LLM-Seq based on GPT-4 overlap significantly on all datasets except Diabetes Progression and NYC Rideshare* (Figure 3(a)). Figures 2 and 3(a) together illustrate that GPT-4 achieves consistently strong feature selection performance regardless of the selection mechanism. Notably, even the simple strategy of querying for an importance score one-feature-at-a-time can be just as effective as those that account for the other features. We overall observe a similar trend in the feature selection paths for GPT-3.5 (Figure 3(b)), which is albeit less pronounced. Meanwhile, for the smaller Llama-2 models, the performances of LLM-Score, LLM-Rank, and LLM-Seq are less consistent across datasets (Figures A3–A5), and no single method necessarily outperforms the others. These results suggest that while all three LLM-based feature selection methods can be similarly effective, their effectiveness varies more significantly across datasets as the model size decreases.
Result 3: Zero-shot prompting with no dataset-specific context and greedy decoding results in strong feature selection performance (Figure 4).
To assess the sensitivity of LLM-Score to prompt design and decoding strategy variations, we evaluate the downstream test performance across six prompt designs (“prompt design” in Section 4) and two decoding strategies (greedy/self-consistency). We use the change in the area under the feature selection paths to measure the impact of each variation. For classification tasks, an increase in the area indicates improved performance, as it suggests that fewer features are needed to achieve high AUROC. Conversely, for regression tasks, a decrease in the area indicates improvement. Thus, we quantify improvement in feature selection performance by computing the % increase for the classification tasks and the % decrease for the regression tasks. Figure 4 shows the average improvement for each (prompt design, decoding strategy) pair—with respect to that achieved with the default prompt and greedy decoding (). For the settings where we add examples, we only consider a one-shot setting given the relatively small number of features in the datasets. We find that no prompting strategy consistently improves feature selection performance across all LLMs, sometimes even degrading it (e.g., GPT-3.5, Llama-2 (13B)). Meanwhile, we find that the largest and most capable GPT-4 is substantially less sensitive to the prompting strategy and generally benefits from additional context, albeit to a limited extent. These observations suggest that zero-shot prompting without dataset-specific context, combined with greedy decoding, is a strong baseline prompting strategy for LLM-based feature selection.
Result 4: LLM-Score exhibits higher correlation with widely used feature importance metrics as model scale increases (Figure 5).
@@
We compute the Kendall’s coefficient (Kendall, 1938) to quantify the agreement in pairwise orderings, where +1/-1 indicates perfect agreement/disagreement. For SHAP, we compute the mean absolute Shapley value for each feature using all test samples, after training an XGBoost model (Chen & Guestrin, 2016). For permutation importance, we measure the average drop in test performance of an -penalized logistic/linear regression model when randomly shuffling the values of each feature 30 times. In Figure 5, we observe that there is no specific notion of importance that LLM-Score consistently aligns to. Meanwhile, as model scale increases, the LLM-generated scores generally exhibit higher correlation with the importance metrics considered (e.g., SHAP, Spearman). An interesting future research direction may be to investigate prompting strategies that increase the alignment of LLM-Score to desired notions of importance. We include additional results on investigating the semantics of LLM-Score in Appendix A.3.7.
@@ 4.2 Evaluation on Large-Scale Datasets
We show that LLM-based feature selection also achieves strong performance on the more complex large-scale, higher dimensional datasets, each with 3000 features after preprocessing. Here, we focus on GPT-4-based LLM-Score with the default prompting setup, given its consistently strong performance on small-scale datasets and scalability to a large number of features. We construct supersets of the Income, Employment, Public Coverage, and Mobility datasets from folktables (Ding et al., 2021) by extracting all features available from the 2018 American Community Survey data for California while removing features that lead to label leakage. We also manually extract three datasets from the MIMIC-IV database (Johnson et al., 2023) for classifying whether an ICU patient was diagnosed with chronic kidney disease (CKD), chronic obstructive pulmonary disease (COPD), and heart failure (HF). Importantly, we note that MIMIC-IV is not publicly available (access requires special credentials via PhysioNet (Goldberger et al., 2000)) and that the datasets we manually derived are not based on existing data preprocessing pipelines (e.g., MIMIC-Extract (Wang et al., 2020), FIDDLE (Tang et al., 2020)). As such, these exact datasets were not part of the LLM pretraining corpora. We provide the remaining dataset details in Appendix B.1.
We evaluate each feature selection method by measuring how the test performance of a downstream prediction model changes when we select the top 10%, 30%, 50%, 70% or 100% of all input concepts. We train the following models for downstream prediction: LightGBM (Ke et al., 2017), MLP, and -penalized logistic regression. For each feature selection method and dataset, we perform model selection via a random search with 40 hyperparameter samples. We average the test AUROC over 5 random seeds, which control the train-validation splits, the hyperparameter samples, and the initialization of model parameters.
@@
the LASSO (Tibshirani, 1996),
@@
minimum redundancy maximum relevance (MRMR; Ding & Peng, 2005),
@@
HSIC-Lasso (Yamada et al., 2014),
@@
Concrete Autoencoder (CAE; Balın et al., 2019),
6.
random feature selection.
Unlike for the small-scale dataset experiments, we exclude forward sequential selection, backward sequential selection, recursive feature elimination (Guyon et al., 2002), and Sequential Attention (Yasuda et al., 2023), given the high computational cost of training separate models when selecting out of candidate features. For LassoNet, the LASSO, and the random feature selection baseline, we set up each method to select features at the concept level (e.g., “ethnicity” instead of the one-hot encoded features “ethnicity_Asian” or “ethnicity_Hispanic”) to match the behavior of LLM-Score (e.g., via group-wise feature sparsity (Yuan & Lin, 2006)). For LassoNet and the LASSO, we compute the regularization paths with warm starts (Friedman et al., 2010) (Figure B1) and use the regularization coefficient that selects about the same number of features as LLM-Score in each evaluation setting. For MRMR, HSIC-Lasso, and CAE, we select exactly the same number of features as LLM-Score in each setting. We use the same setup for HSIC-Lasso as in Section 4.1 and additionally subsample down to 60000 training examples if the training set size exceeds it to avoid out-of-memory errors (Employment and Public Coverage; Appendix B.1). For CAE, we adopt the setup in Section 4.1 but increase the hidden layer widths for the MLP decoder from 256 to 512. We provide the remaining details in Appendix B.2.
@@
Figure 6(a) shows the downstream test AUROC, averaged over all of the folktables (left) and MIMIC-IV (right) datasets, when (i) selecting the top 30% of features according to each baseline and LLM-Score based on GPT-4 and (ii) using LightGBM as the downstream prediction model. Here, we show the performance when selecting the top 30% of features, as the test performance roughly saturates to that of using all features when selecting more than 30% of features for all feature selection methods (Figure B2(a)). We observe that GPT-4-based LLM-Score overall performs the best on the folktables datasets, and performs on par with HSIC-Lasso and significantly better than LassoNet and the random feature selection baseline on the MIMIC-IV datasets. In Figures 6(b–c), we show the full feature selection paths for the best-performing baseline (MRMR), LLM-Score, and random feature selection on the Income and Employment datasets from folktables and the CKD and COPD datasets from MIMIC-IV. We observe that for most of these datasets, only a small subset of features are highly predictive of the target outcome, as indicated by the high test AUROC of the best-performing baseline and the low test AUROC of the random feature selection baseline at the 10% and 30% marks. We find that LLM-Score is effective at selecting these highly predictive features, with performance rivaling that of the best-performing data-driven baseline and substantially better than random when selecting the top 10% and 30% of features. Notably, the strong performance on the MIMIC-IV datasets suggests that LLM-Score can be effective even in highly specialized domains like healthcare where, without access to sufficient training data, substantial domain knowledge would be required for feature selection (Section B.1.2). We include the full results for all downstream prediction models on all datasets in Figure B2 in Appendix B.3.
@@ 5 Discussion and Conclusion
In this work, we demonstrated that LLMs are capable of performing feature selection for supervised learning tasks, even without access to the downstream training data. We proposed three different approaches to prompting LLMs for feature selection—LLM-Score, LLM-Rank, and LLM-Seq (Section 3). We found that with sufficient LLM scale, even zero-shot prompting an LLM to select features can result in strong downstream predictive performance, often competitive with data-driven methods such as the LASSO (Tibshirani, 1996) (Results 1 & 3, Section 4.1). For the latest models such as GPT-4, we found that all three LLM-driven methods achieve similarly strong performance and that even the simple strategy of querying for an importance score one-feature-at-a-time (LLM-Score) can be as effective as those that account for other available features (Result 2, Section 4.1; Section 4.2). In particular, while the LLM-generated scores do not necessarily align to a specific notion of feature importance (e.g., Fisher score, mutual information), we found that their rank correlations tend to be higher as model scale increases (Result 4, Section 4.1), suggesting that the most capable LLMs are able to translate the real-world relationships encoded in their parameters into statistically meaningful scores. Our findings suggest that LLMs may be useful not only for selecting the best features post data collection, but also for deciding what features to collect in the first place. This could potentially benefit practitioners in domains like healthcare and the social sciences, where collecting high-quality data can be expensive and time-consuming.
Limitations.
First, while we have demonstrated the effectiveness of LLM-based feature selection methods in various domains (e.g., finance, healthcare, criminal justice), it is possible that they show limited performance in extremely specialized or rare domains. Second, as LLM-based feature selection methods rely on textual semantics (e.g., column names in tabular datasets) to identify predictive features, they may not be effective on (i) datasets without sufficient text annotations (Appendix A.3.6) or (ii) datasets where the statistical relationships between features and targets do not align with the semantic associations inferred by the LLM (due to e.g., selection bias). Third, LLMs may exhibit undesirable biases inherited from their pretraining data (Gallegos et al., 2024), which can result in downstream performance disparities across data subpopulations. In the data-driven setting, it is possible to mitigate such issues by selecting features for each subpopulation independently or modifying the training objective to account for group fairness (Zemel et al., 2013; Sagawa et al., 2020; Izmailov et al., 2022). However, it is not immediately obvious how to incorporate similar notions of group fairness into LLM-based feature selection methods, especially for closed-source LLMs which can only be accessed via prompting. As such, combining LLM-driven feature selection with data-driven methods or using it in a human-in-the-loop setup may be a more reliable approach for mitigating bias concerns, especially in safety-critical domains. We leave an in-depth investigation of methods for incorporating group fairness to LLM-based feature selection for future work.
@@ Acknowledgements
We gratefully acknowledge DARPA (FA8750-23-2-1015), ONR (N00014-23-1-2368), NSF (IIS2211955), UPMC, Highmark Health, Abridge, Ford Research, Mozilla, the PwC Center, Amazon AI, JP Morgan Chase, the Block Center, the Center for Machine Learning and Health, and the CMU Software Engineering Institute (SEI) via Department of Defense contract FA8702-15-D-0002, for their generous support of our research. We also thank Michael Oberst, Goutham Rajendran, Bingbin Liu, and Che-Ping Tsai for helpful feedback and comments on earlier versions of this work.
@@ A.1 Datasets
For each dataset, we randomly shuffle and take a 80–20 train–test split. We then take a 5-fold split of the training set for cross-validation, where the 5-fold splits vary across the random seeds (=[1,2,3,4,5]) used throughout the experiments. The test set remains fixed and does not vary with the random seed used. For classification datasets, we always take a stratified split to preserve the label proportions across the train, validation, and test sets. We standardize all of the numerical features to have zero mean and unit variance, and one-hot encode all categorical features. Below, we provide the remaining details for each dataset.
@@
Credit-G (Hofmann, 1994) is a UCI dataset555https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data, where the goal is to predict whether a client at a bank carries high credit risk, given a set of attributes about the client (e.g., credit history, savings account status). The dataset contains 1000 samples and 61 features after preprocessing (20 concepts), with 700 positive samples and 300 negative samples.
@@
Bank (Moro et al., 2012) is a UCI dataset666https://archive.ics.uci.edu/dataset/222/bank+marketing, where the goal is to predict whether a client at a bank will subscribe to a term deposit, given data collected from a telemarketing campaign at a Portuguese banking institution from 2008 to 2013. The dataset contains 45211 samples and 51 features after preprocessing (16 concepts), with 5289 positive samples and 39922 negative samples.
@@
Give Me Some Credit is a Kaggle dataset777https://www.kaggle.com/c/GiveMeSomeCredit, where the goal is to predict whether an individual is likely to experience significant financial distress/delinquency within the next two years, given information about the individual’s financial status (e.g., debt ratio, monthly income). The dataset contains 120269 samples and 10 features after preprocessing (10 concepts), with 8357 positive samples and 111912 negative samples.
@@
COMPAS Recidivism is a dataset collected from a 2016 study888https://www.propublica.org/article/how-we-analyzed-the-compas-recidivism-algorithm on the racial biases present in the Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) algorithm (Larson et al., 2016). The goal is to predict whether a criminal defendant carries high risk of recidivism, given their criminal history and demographic attributes. The dataset contains 6172 samples and 23 features after preprocessing (14 concepts), with 2751 positive samples and 3421 negative samples.
@@
Pima Indians Diabetes (Smith et al., 1988) is a Kaggle dataset999https://www.kaggle.com/datasets/uciml/pima-indians-diabetes-database, where the goal is to predict whether a female adult patient of Pima Indian heritage has diabetes, given a set of clinical measurements and demographics. The dataset contains 768 samples and 8 features after preprocessing (8 concepts), with 268 positive samples and 500 negative samples.
@@
AUS Cars* is a Kaggle dataset101010https://www.kaggle.com/datasets/nelgiriyewithana/australian-vehicle-prices published in 2023, where the goal is to predict whether the price of a car in Australia is above $30000, given a set of attributes about the car (e.g., year of manufacture, number of doors in the car, fuel consumption rate). The dataset contains 14283 samples and 33 features after preprocessing (11 concepts), with 6396 positive samples and 7887 negative samples.
@@
YouTube* is a Kaggle dataset111111https://www.kaggle.com/datasets/nelgiriyewithana/global-youtube-statistics-2023 published in 2023, where the goal is to predict whether YouTube channel has more than 20 million subscribers, given a set of attributes about the channel (e.g., total number of videos uploaded on the channel, date when the channel was created). The dataset contains 588 samples and 72 features after preprocessing (22 concepts), with 268 positive samples and 320 negative samples.
@@
CA Housing (Pace & Barry, 1997) is a StatLib dataset121212https://www.dcc.fc.up.pt/~ltorgo/Regression/cal_housing.html, where the goal is to predict the median housing price of a US Census block group in California, given data collected from the 1990 US Census. A block group, typically a population of 600 to 3000 people, is the smallest geographical unit for which the US Census Bureau publishes sample data. The dataset contains 20640 samples and 8 features after preprocessing (8 concepts).
@@
Diabetes Progression is a dataset131313https://www4.stat.ncsu.edu/~boos/var.select/diabetes.html used in a study by Efron et al. (2004), where the goal is to predict the disease progression level in diabetic patients, given their baseline blood serum measurements from the previous year and demographic information. The dataset contains 442 samples and 10 features after preprocessing (10 concepts).
@@
Wine Quality (Cortez et al., 2009) is a UCI dataset141414https://archive.ics.uci.edu/dataset/186/wine+quality collected from red and white vinho verde wine samples from northern Portugal, where the goal is to predict whether a wine is high or low quality, given its various physicochemical measurements (e.g., acidity, density). The dataset contains 6497 samples and 11 features after preprocessing (11 concepts).
@@
Miami Housing (Grinsztajn et al., 2022) is an OpenML dataset151515https://www.openml.org/search?type=data&sort=runs&id=43093&status=active, where the goal is to predict the selling price of a house in Miami in 2016, given structural (e.g., area, structure quality) and geographic (e.g., longitude, latitude) information about each house. The dataset contains 13932 samples and 15 features after preprocessing (15 concepts).
@@
Used Cars is a Kaggle dataset161616https://www.kaggle.com/datasets/vijayaadithyanvg/car-price-predictionused-cars, where the goal is to predict the selling price of a used car, given a set of attributes about each car (e.g., age, fuel type, number of previous owners). The dataset contains 301 samples and 11 features after preprocessing (7 concepts).
@@
NBA* is a Kaggle dataset171717https://www.kaggle.com/datasets/bryanchungweather/nba-player-stats-dataset-for-the-2023-2024 published in 2023, where the goal is to predict the number of points per game for an NBA basketball player, given a set of attributes and statistics (from the 2023–2024 NBA season) about the player (e.g., blocks per game, position, team). The dataset contains 388 samples and 61 features after preprocessing (27 concepts).
@@
NYC Rideshare* is a Kaggle dataset181818https://www.kaggle.com/datasets/aaronweymouth/nyc-rideshare-raw-data published in 2023, where the goal is to predict the total pay given to a rideshare-app driver in NYC after a trip, given information about the trip (e.g., time elapsed from ride request to dropoff, total distance of the trip). The original dataset on Kaggle contains some data from 2022, so we only extract the data rideshare data from 2023 for preprocessing and evaluation. The dataset contains 5000 samples and 19 features after preprocessing (15 concepts).